OpenStudy (anonymous):

In a 45°-45°- 90° triangle, if the legs each have a length of 4, then what is the length of the hypotenuse?

OpenStudy (anonymous):

pytthagorean theorem 4^2 + 4^2 = c^2 16 + 16 = c^2 32 = c^2 c = sqrt(32) <- hypotenuse

OpenStudy (anonymous):

multiple choices are a. sqrt 2 b. 4 sqrt 2 c. 4 d. 4sqrt 3

OpenStudy (anonymous):

if you simpllfied sqrt(32) it becomes 2sqrt(8) =4sqrt(2) so b

OpenStudy (anonymous):

Just remember that the hypotenuse is always the leg √ 2 that's a shortcut behind all of the math! ex: if your leg was 5 the your hypotenuse would be 5√ 2

OpenStudy (anonymous):

This method is used when dealing with an isosceles triangle/ 45 45 90 degree triangle.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

no problem!